PITMinutes 07-17-2020

From CASA Wiki
Jump to: navigation, search


K. Baggett, J. Benesch, M. Bickley, T. Carlino, G. Ciovati, J. Creel, E. Daly, K. Davis, M. Drury, C. Hovater, N. Huque, F. Humphrey, T. Ganey, C. Ginsburg, J. Kent, A. Kimber, L. King, G. Krafft, F. Marhauser, R. Michaud, C. Mounts, R. Nelson, W. Oren, A. Palczewski., T. Plawski, T. Powers, J. Preble, C. Reece, T. Reilly, A. Seryi, A. Solopova, M. Spata, and P. Vasilauskis

Agenda and Notes

1. Operations Status and Summary/Review (Bickley/Drury)

Results of 1L25 FE measurements

Summary of Field Emission Measurements

2. Energy Reach (Mounts)

RF/SRF issues and current energy reach

14 cavities hard off in North linac 4 cavities hard off in South linac

Lem to 1031/linac has very high trip rate at moment.

Non-symmetrical running leads to loss of polarization in one of Hall A or Hall C

3. General discussion

Action items:[1] Continue analysis of two hour pressure spikes and check module stability after problem solved. [2] Increase Q_ext to allow higher operating gradient in 4 cavities [3] Turn on 1L08, in particular cavity three at low gradient. Hold field at low set point to avoid uncontrolled gradient in cavity (cannot be detuned away from fundamental frequency)